# Dynamical critical behavior on the Nishimori point of frustrated Ising models

### v1

@@ III.4 Multifractality of moments
Figure 9: Dynamical length scale $\xi_{q}(t)$ vs time $t$ for different odd values of $q$ (see the key). The dashed line represents the power law, $\xi_{q}(t)\sim t^{1/z_{\rm c}}$, with $z_{\rm c}\simeq 6$.
@@ IV Summary and discussion
In this paper, we investigated different aspects of the critical dynamics (generated by single spin flip non-conserving order parameter) at the NP in the two-dimensional frustrated Ising models. We calculated the three dynamical critical exponents — $z_{\rm c}$, $\lambda_{\rm c}$, and $\theta_{\rm c}$. To establish universality of these exponents, we defined the model on different lattice geometries (SL, TL, and HL) with the quenched bond variables drawn from bimodal (±$(\pm J)$ and Gaussian distributions. We numerically found that after a quench from the high temperature phase to the NP, different dynamical exponents attain their universal values. Therefore, the effects of frustration which significantly modify the critical properties of a pure system leave the universality in the critical exponents unchanged. This also validates the RG picture of the critical phenomenon in the presence of frustration, i.e., the microscopic details do not affect the large scale physics.
We found that due to systematic errors the value of the exponent $z_{\rm c}$ obtained from power-law fits fluctuates between $6.0-6.5$ for different geometries and bond distributions. The effective exponent $z_{\rm eff}$, defined in Eq. (16), also endorses the same order of error in $z_{\rm c}$. Notice that the focus of the present work was to establish the universality in the exponents, not the precise determination of the exponent values. In our recent study on the SL with ±$\pm J$ disorder Agrawal et al. (2023a) (also see Ref. Ozeki and Ito (1998)), we observed $z_{\rm c}\sim 6$ at late times ($t\sim 10^{7}$) and large system $(L=1024)$. Whereas, in the present study, timescales up to $t\sim 10^{6}$ on $L=512$ are accessed. Given that, also invoking Occam’s razor, a universal exponent $z_{\rm c}\sim 6$ is expected in the larger simulations on effectively larger system sizes.

### v2

@@ III.4 Multifractality of moments
Figure 9: Dynamical length scale $\xi_{q}(t)$ vs time $t$ for different odd values of $q$ (see the key). The dashed line represents the power law, $\xi_{q}(t)\sim t^{1/z_{\rm c}}$, with $z_{\rm c}=6$.
@@ IV Summary and discussion
In this paper, we investigated different aspects of the critical dynamics (generated by single spin flip non-conserving order parameter) at the NP in the two-dimensional frustrated Ising models. To the best of our understanding, this is the first study in which the dynamical properties of the multicritical NP are investigated in detail. We calculated the three dynamical critical exponents — $z_{\rm c}$, $\lambda_{\rm c}$, and $\theta_{\rm c}$. To establish universality of these exponents, we defined the model on different lattice geometries (SL, TL, and HL) with the quenched bond variables drawn from bimodal (±$(\pm J)$ and Gaussian distributions. We numerically found that after a quench from the high temperature phase to the NP, different dynamical exponents attain their universal values, which confirms that the effects of frustration leave the universality in the dynamical exponents unchanged. This also further consolidates the RG picture of dynamical critical phenomenon in the presence of frustration.
We found that due to systematic errors the value of the exponent $z_{\rm c}$ obtained from power-law fits fluctuates between $6.0-6.6$ for different geometries and bond distributions. The effective exponent $z_{\rm eff}$, defined in Eq. (16), also endorses the same order of error in $z_{\rm c}$. Notice that the focus of the present work was to establish the universality in the exponents, not the precise determination of the exponent values. In our recent study on the SL with ±$\pm J$ disorder Agrawal et al. (2023a) (also see Ref. Ozeki and Ito (1998)), we observed $z_{\rm c}\sim 6$ at late times ($t\sim 10^{7}$) and large system $(L=1024)$. Whereas, in the present study, timescales up to $t\sim 10^{6}$ on $L=512$ are accessed. Given that, also invoking Occam’s razor, a universal exponent $z_{\rm c}\sim 6$ is expected in the larger simulations on effectively larger system sizes.